# Quadratic Mode Couplings in Rotating Black Holes and Their Detectability

### v2

@@ Abstract
Quadratic quasinormal modes encode fundamental properties of black hole spacetimes. They are also one of the key ingredients of nonlinearities of General Relativity in the ringdown stage of binary black hole coalescence. In this work, we classify all possible quadratic coupling channels of quasinormal modes for a generic Kerr black hole, and use a frequency-domain pseudospectral code with hyperboloidal slicing to calculate these couplings. After accounting for all the channels in systems with reflection symmetry, our results become consistent with those extracted from numerical simulations and time-domain fits. We explore several observationally relevant quadratic modes and their detectability with third-generation ground-based detectors like Cosmic Explorer, as well as the space-borne detector LISA (Laser Interferometer Space Antenna). We find that several quadratic modes are detectable by both detectors. In addition, we make a theoretical prediction for how the signal-to-noise ratio of a quadratic quasinormal mode depends on the properties of progenitor binaries, offering insight into selecting golden events in future ringdown data analyses.
@@ I Introduction
Among various BH non-linearities [10, 11, 12, 13, 14, 7, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], quadratic quasinormal modes (QQNMs) stand out due to their conceptual simplicity and their potential observability using current ringdown fitting techniques [27, 28, 29, 30, 31, 32, 33]. QQNMs have been identified by fitting numerical time-domain simulation data at infinity [10, 11, 12, 13, 14, 7] and on the BH horizon [17]. However, a complete and accurate characterization of the modes relies on frequency domain methods, which is the focus of this work.
Along with an earlier work [18] (hereafter Paper I), we develop two independent frequency-domain methods to handle QQNMs for generic Kerr BHs. In Paper I, we analytically reconstructed the metric and solved the frequency-domain second-order Campanelli–Lousto-Teukolsky equation [34] along a complex contour. Alternatively, in this paper, we adopt the reduced second-order Teukolsky equation [35, 36] and use the pseudo-spectral method to numerically assemble its source term. Subsequently, we solve the equation spectrally in hyperboloidal slicing [37, 38, 39]. Both methods generate consistent results for the coupling coefficient of the prograde mode $\ell=m=2,n=0$ for BH spins $\chi\in[0,0.95]$. Furthermore, we extend our discussions in Paper I by studying more observationally relevant QQNMs, and by providing a detailed classification of all possible mode-coupling channels. After examining mode coupling through different channels and symmetry conditions in (some of the) previous numerical simulations, we find consistency with the time-domain fits for mode amplitudes.
Leveraging the recent ringdown surrogate model [^1] See also [9]. [8], we obtain theoretical predictions for how the amplitudes of quadratic modes depend on the properties of progenitor binaries. Assuming the sensitivities of Cosmic Explorer (CE) and Laser Interferometer Space Antenna (LISA) [40, 41], we carry out a parameter survey to forecast the SNRs of several leading quadratic modes in non-precessing binaries, complementing the recent study by Yi et al. [42]. Our findings indicate that several QQNMs are within the detectable range of both detectors.
Throughout this work we adopt the geometric unit by setting $G=c=1$. Complex conjugates are represented by overlines. We label a QNM with four indices $(l,m,n,p)$, where $p={\rm sgn}({\rm Re}(\omega))$. QNMs come in pairs. Their frequency values are related via
@@
which indicates that the modes with $(m,p)$ and $(-m,-p)$ (and the same $l,n$) are mirrors of each other.
@@ II QQNMs with hyperboloidal slicing
Our numerical implementation uses the pseudo-spectral method, where all quantities are expanded in Chebyshev polynomials radially and spin-weighted spherical harmonics in the angular directions. Multiplications are performed using fast transforms from the spectral basis to a grid. All results in this paper have radial expansion truncated at degree $n_{\text{max}}=63$, the spherical harmonic expansion truncating at $\ell_{\text{max}}=31$ and with 256 bit floating point numbers from the MultiFloats library [45].
To obtain the source term $S_{4}^{(2)}$, we take a different approach from the analytic treatment used in Paper I—here, the source is computed numerically from the Einstein Tensor. To be specific, we first consider the second-order perturbation of the metric
@@
where $G_{ab}^{(1)}$ and $2G_{ab}^{(2)}$ are the first and second functional derivatives of the Einstein Tensor with respect to the metric. To first order, we take $h_{ab}^{(1)}$ to be a sum of QNMs. We solve for $h_{ab}^{(1)}$ by numerically computing the Hertz potential for QNMs in the Outgoing Radiation Gauge (ORG), using the separation technique described in [44]. Note that for a QNM with frequency $\omega$, $h_{ab}^{(1)}$ will have both frequencies $\omega$ and $-\bar{\omega}$, which comes from the complex conjugations necessary to make $h_{ab}^{(1)}$ real.
@@
thus $h_{ab}^{(2)}$ satisfies the linearized Einstein equations with $-G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]/8\pi$ acting as the effective stress-energy tensor at this order. After computing $G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]$ from the linearized metric $h_{ab}^{(1)}$, we convert it to the source $S_{4}^{(2)}$ for $\Psi_{4}^{(2)}[h_{ab}^{(2)}]$ in Eq. (2). Note that while Eq. (5) only provides the source for the part of $\Psi_{4}$ that depends linearly on $h_{ab}^{(2)}$, in ORG quadratic contribution to $\Psi_{4}$ from $h_{ab}^{(1)}$ vanishes.
@@ III Classification of mode couplings
Quadratic coupling channels can be viewed in two ways. The first perspective traces the relationship from linear parent modes to quadratic daughter modes. Consider two linear QNMs $(\ell_{1},m_{1},n_{1},p_{1})$ and $(\ell_{2},m_{2},n_{2},p_{2})$, and amplitudes $A_{1}$ and $A_{2}$. As discussed above, the linearized metric $h_{ab}^{(1)}$ has four frequency components $\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}$, $-\bar{\omega}_{\ell_{1}m_{1}n_{1}}^{p_{1}}$, $\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}}$, and $-\bar{\omega}_{\ell_{2}m_{2}n_{2}}^{p_{2}}$, with the amplitudes $A_{1}$, $\bar{A}_{1}$, $A_{2}$ and $\bar{A}_{2}$. Consequently, the source term in Eq. (5) yields four quadratic frequencies:
@@
with source terms proportional to the products $A_{1}A_{2}$, $A_{1}\bar{A}_{2}$, $\bar{A}_{1}A_{2}$ and $\bar{A}_{1}\bar{A}_{2}$, respectively. This gives rise to four distinct channels through which a pair of linear QNMs can generate QQNMs at different frequencies.
The first perspective, which maps two parent modes to four daughter modes, is often conceptually inconvenient. Instead, it is more intuitive to adopt the opposite viewpoint: given a QQNM with a certain frequency, what are its parent modes? To formalize this, we will use the following notation to denote a QQNM:
@@
With this notation, several key observations can be made. First, the QQNM appears only in the $(m_{1}+m_{2})$ GW harmonics due to the azimuthal angular selection rule. Second, the notation exhausts all possible QQNMs. Third, different pairs of modes (up to exchange symmetry $1\leftrightarrow 2$) generally produce distinct frequencies. An exception occurs for Schwarzschild BHs, where spherical symmetry makes QNM frequencies independent of $m$. Nonetheless, different $(m_{1},m_{2})$ pairs indicate distinct coupling channels. The total QQNM amplitude in a GW harmonic is the sum of all possible pairs $(m_{1},m_{2})$ with $m_{1}+m_{2}$ fixed.
The QQNM in Eq. (6) is generated by four linear QNMs:
@@
with respective amplitudes $A_{1}$, $A_{-1}$, $A_{2}$, and $A_{-2}$. It is worth noting that while the frequency values of $(\ell_{1},m_{1},n_{1},p_{1})$ and $(\ell_{1},-m_{1},n_{1},-p_{1})$ are related via Eq. (1), their amplitudes $A_{1}$ and $A_{-1}$ are independent, except in the case of nonprecessing systems, where because of the reflection symmetry we have
@@
The four linear QNMs are coupled through four channels
@@
Below we will refer to them as +++++ +, ++-+ -, +-+- +, and --- -, respectively. For example, the (+)(+-)( + - ) channel corresponds to the interaction where the linear mode $(\ell_{2},-m_{2},n_{2},-p_{2})$ generates a frequency component $-\bar{\omega}_{\ell_{2}-m_{2}n_{2}}^{-p_{2}}$ in the linearized metric, which is numerically equal to $\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}}$ due to Eq. (1). This component then interacts with the frequency $\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}$ generated by the first linear mode $(\ell_{1},m_{1},n_{1},p_{1})$. Denoting the amplitude of the QQNM as $A_{Q}$, we have
@@
where the coefficients $R$’s are fully determined by the dimensionless spin of a Kerr BH. Here we use spin-weighted spherical/spheroidal harmonics as the mode basis to facilitate direct comparison with previous results, such as those from numerical relativity and from Paper I. Alternatively, modes can also be classified by parity, see the Supplementary Material for more details.
@@
The structure of $S_{4}^{(2)}$ explicitly illustrates how two linear parent modes couple. The first parent mode contributes the three Weyl scalars $\Psi_{4}^{(1)}$, $\Psi_{3}^{(1)}$, and $\Psi_{2}^{(1)}$, which are then multiplied by the terms contributed by the second parent mode (indicated in parentheses). For a given linear QNM $(l,m,n,p)$, the expressions for $\Psi_{4}^{(1)}$, $\Psi_{3}^{(1)}$, and $\Psi_{2}^{(1)}$ were analytically derived in Paper I [18], see Eqs. (38), (42), and (46) therein. An intriguing feature emerges: all three Weyl scalars contain only a single frequency component, $\omega_{lmn}^{p}$, with no contribution from their mirrored counterpart, $-\bar{\omega}_{lmn}^{p}$. As a result, the source in Eq. (12) contains no terms resulting from the product of two mirrored frequency components.
@@ IV Results and Comparisons
The QQNM $\begin{pmatrix}220+\\ 220+\end{pmatrix}$ in the $l=m=4$ harmonic has been extensively discussed in the literature [12, 10, 11, 13, 14, 19, 20, 21, 18]. In particular, in Paper I, we found that the computed amplitude ratio, which is essentially $R_{++}$ in Eq. (10), was smaller than those extracted from numerical relativity [13, 14]. We speculated that the difference was due to parity freedom in the linear parent modes (see Sec. VI. B in [18]). This was later confirmed in [19, 20, 21] for Schwarzschild BHs. In this work, after classifying all mode-coupling channels, we revisit the discrepancy for Kerr BHs.
@@
Besides $\begin{pmatrix}220+\\ 220+\end{pmatrix}$, we also study five more QQNMs, including $\begin{pmatrix}330+\\ 220+\end{pmatrix}$, $\begin{pmatrix}440+\\ 2\text{-}20-\end{pmatrix}$, $\begin{pmatrix}220+\\ 200+\end{pmatrix}$, $\begin{pmatrix}220+\\ 200-\end{pmatrix}$, and $\begin{pmatrix}330+\\ 2\text{-}20-\end{pmatrix}$. In particular, we use polynomials to fit their $R$ for different channels as functions of $\chi$:
@@
V Forecast on detectability
Using the total amplitude ratio $R^{\rm total}$ from Eq. (11), we can estimate the SNR of various QQNMs emitted by nonprecessing binaries. To be specific, for a given binary system, we first adopt the recent surrogate model postmerger [8] to compute linear mode amplitudes, along with the mass and spin of the remnant BH. We then use our fitting formula for $R^{\rm total}$ in Eq. (13) to estimate the amplitude of the corresponding QQNMs. Finally, we follow the method outlined in [46, 47, 42] to compute their SNRs.
In the following, we focus on a 40 km CE [48] and LISA [49]. For simplicity, we average the SNRs over all angles, which yields the orthogonal condition between two antenna patterns $F_{+,\times}$ for CE:
@@ V Forecast on detectability
For LISA, there is an additional factor of $3/2$. As pointed out in Paper I, a QQNM is present in different $\ell$ harmonics. Its total SNR is given by given by
@@
(a) $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$
(b) $\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$
@@
Figure 2: The SNR dependence of two QQNMs, $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$, and $\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$, on progenitor binaries’ mass ratio $q$ and effective spin $\chi_{\rm eff}$. The systems have a total mass of $100M_{\odot}$ and a luminosity distance of 410 Mpc. CE sensitivity is used.
@@
In this paper, we adopt a conservative estimate by choosing the starting time to be $10M_{\rm tot}$ after the peak, where $M_{\rm tot}$ is the total mass of the binary system. Giesler et al. [50] demonstrate stable QNM fits at earlier times. Selecting an earlier starting time can moderately increase the SNRs, since QQNMs decay faster than linear fundamental modes.
The SNR of $\begin{pmatrix}2,&2,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$ was investigated in [42]. In the Supplementary Material, we present our results. A major difference between us is in the SNR’s dependence on the mass ratio $q$. In their Fig. 6, the SNR peaks at $q\gtrsim 2$, while our more accurate calculation of the amplitude ratio shows a monotonic decrease with $q$. Therefore, the golden binary for detecting the QQNM is an equal-mass system with maximally positive spins.
The QQNM $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$ was also identified in numerical relativity waveforms. In Fig. 2a, we present its SNR dependence on the mass ratio $q$ and the effective inspiral spin $\chi_{\rm eff}$. For simplicity, we assume both component spins $\chi_{1,2}$ are equal, namely $\chi_{1}=\chi_{2}=\chi_{\rm eff}$. The optimal binary configuration is $q\sim 2.75$ with maximally positive $\chi_{\rm eff}$. By setting the total binary mass $M_{\rm tot}$ to $100M_{\odot}$, and the luminosity distance $D_{L}$ to 410 Mpc, the SNR detected by CE can exceed 8 for some parameter space. With LISA, the SNR dependence on $q$ and $\chi_{\rm eff}$ remains the same. Assuming $M_{\rm tot}=8\times 10^{6}M_{\odot}$ and a redshift of 1, we find its SNR ranges from 16 to 653. The corresponding plot is available in the Supplementary Material.
The QQNM $\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$ is weaker. Comparing to $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$, this QQNM has a smaller frequency and appears in the $(\ell\geq 2,m=1)$ harmonics. As shown in Fig. 2b, its SNR still peaks at $q\sim 2.75$ with maximally positive $\chi_{\rm eff}$. Using the same binaries, the SNR ranges from $0.1$ to $10.6$ for CE, and from $0.5$ to $59.8$ for LISA (see the LISA plot in the Supplementary Material).
Here we focus on two typical binary systems for CE and LISA by fixing their $M_{\rm tot}$ and $D_{L}$. The impact of $M_{\rm tot}$ on SNR is two-fold: (1) it scales the overall strain linearly; and (2) it sets the central frequency of a QQNM. In the Supplementary Material, we show how the SNR of $\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$ changes with $M_{\rm tot}$. For CE, SNR increases monotonically from 0.6 to 30 over $M_{\rm tot}\in[20,200]M_{\odot}$, while for LISA, it reaches a maximum for $M_{\rm tot}=5.3\times 10^{6}M_{\odot}$, with a value of 64.

### v3

@@ Abstract
Quadratic quasinormal modes encode fundamental properties of black hole spacetimes. They are also one of the key ingredients of nonlinearities of General Relativity in the ringdown stage of binary black hole coalescence. In this work, we classify all possible quadratic coupling channels of quasinormal modes for a generic Kerr black hole, and use a frequency-domain pseudospectral code with hyperboloidal slicing to calculate these couplings. After accounting for all the channels in systems with reflection symmetry, our results become consistent with those extracted from numerical simulations and time-domain fits. This agreement provides a compelling example demonstrating the success of black hole second-order perturbation theory. We also explore potential applications of our calculations in future ringdown data analysis by carrying out a detectability survey for various quadratic modes. We find that a few of them are observationally relevant for third-generation ground-based detectors like Cosmic Explorer, as well as the space-borne detector LISA (Laser Interferometer Space Antenna).
@@ I Introduction
Among various BH nonlinearities [10, 11, 12, 13, 14, 7, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], quadratic quasinormal modes (QQNMs) stand out due to their conceptual simplicity and their potential observability using current ringdown fitting techniques [27, 28, 29, 30, 31, 32, 33, 34]. QQNMs have been identified by fitting numerical time-domain simulation data at infinity [10, 11, 12, 13, 14, 7] and on the BH horizon [17]. However, a complete and accurate characterization of the modes relies on frequency domain methods, which is the focus of this work.
Along with an earlier work [18] (hereafter Paper I), we develop two independent frequency-domain methods to handle QQNMs for generic Kerr BHs. In Paper I, we analytically reconstructed the metric and solved the frequency-domain second-order Campanelli–Lousto-Teukolsky equation [35] along a complex contour. Alternatively, in this paper, we adopt the reduced second-order Teukolsky equation [36, 37] and use the pseudo-spectral method to numerically assemble its source term. Subsequently, we solve the equation spectrally in hyperboloidal slicing [38, 39, 40]. Both methods generate consistent results for the coupling coefficient of the prograde mode $\ell=m=2,n=0$ for BH spins $\chi\in[0,0.95]$. Furthermore, we extend our discussions in Paper I by providing a comprehensive classification of all possible mode-coupling channels that contribute to the same frequency of quadratic modes. This classification allows us to resolve the disagreement between our previous QQNM calculations in Paper I and numerical simulations.
Leveraging the recent ringdown surrogate model [^1] See also [9]. [8], we discuss potential applications of our calculations in future ringdown data analysis. As an example, we carry out a parameter survey to forecast the SNRs of several quadratic modes measured by Cosmic Explorer (CE) and Laser Interferometer Space Antenna (LISA) [41, 42], complementing the recent study by Yi et al. [43]. Our findings indicate that multiple QQNMs are relevant for future ringdown analysis.
Throughout this work we adopt the geometric unit by setting $G=c=1$. Complex conjugates are represented by overlines. We use $(l,m,n,p)$ to stand for a QNM, where $p={\rm sgn}({\rm Re}(\omega))$; and $\omega$ to denote frequency values. Due to the reflection symmetry of Kerr, QNMs come in pairs with frequency values related via
@@
where modes with $(m,p)$ and $(-m,-p)$ (and the same $l,n$) are mirrors of each other.
@@ II QQNMs with hyperboloidal slicing
To obtain the source term $S_{4}^{(2)}$, we take a different approach from the analytic treatment used in Paper I—here, the source is computed numerically from the Einstein Tensor. Specifically, we consider the second-order perturbation of the metric
@@
where $G_{ab}^{(1)}$ and $2G_{ab}^{(2)}$ are the first and second functional derivatives of the Einstein Tensor with respect to the metric. To first order, we take the GW strain to be a sum of QNMs. We solve for $h_{ab}^{(1)}$ by numerically computing the Hertz potential for QNMs in the Outgoing Radiation Gauge (ORG), using the separation technique described in [45]. Note that for a QNM with frequency $\omega$, $h_{ab}^{(1)}$ will have both frequencies $\omega$ and $-\bar{\omega}$, which comes from the complex conjugations necessary to make $h_{ab}^{(1)}$ real.
@@
thus $h_{ab}^{(2)}$ satisfies the linearized Einstein equations with $-G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]/8\pi$ acting as the effective stress-energy tensor at this order. After computing $G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]$ from the linearized metric $h_{ab}^{(1)}$, we convert it to the source $S_{4}^{(2)}$ for $\Psi_{4}^{(2)}[h_{ab}^{(2)}]$ in Eq. (2). Note that while Eq. (5) only provides the source for the part of $\Psi_{4}$ that depends linearly on $h_{ab}^{(2)}$, in ORG quadratic contribution to $\Psi_{4}^{(2)}$ from $h_{ab}^{(1)}$ vanishes.
@@ III Classification of mode couplings
Quadratic coupling channels can be viewed in two perspectives. The first traces the relationship from linear parent modes to quadratic daughter modes. Consider two linear QNMs $(\ell_{1},m_{1},n_{1},p_{1})$ and $(\ell_{2},m_{2},n_{2},p_{2})$, with strain amplitudes $A_{1}$ and $A_{2}$. The corresponding linearized metric $h_{ab}^{(1)}$ has four frequency components $\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}$, $-\bar{\omega}_{\ell_{1}m_{1}n_{1}}^{p_{1}}$, $\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}}$, and $-\bar{\omega}_{\ell_{2}m_{2}n_{2}}^{p_{2}}$, with the amplitudes $A_{1}$, $\bar{A}_{1}$, $A_{2}$ and $\bar{A}_{2}$, generated by the frequency pairs of the linear modes $(\ell_{1},m_{1},n_{1},p_{1})$ and $(\ell_{2},m_{2},n_{2},p_{2})$. Consequently, the source term in Eq. (5) yields four quadratic frequencies:
@@
with source terms proportional to the products $A_{1}A_{2}$, $A_{1}\bar{A}_{2}$, $\bar{A}_{1}A_{2}$ and $\bar{A}_{1}\bar{A}_{2}$, respectively. This gives rise to four distinct channels through which two linear QNMs can generate QQNMs at different frequencies.
The first perspective, which maps two parent modes to four daughter modes, is often practically inconvenient. Instead, it is more intuitive to adopt the opposite viewpoint: given a certain QQNM frequency, what are its parent modes? To formalize this, we will use the following notation to denote a QQNM frequency component:
@@
This notation represents the set of all QQNMs with that frequency, including various channels, and is not restricted to fixed daughter spherical/spheroidal harmonic. With this notation, several key observations can be made. First, QQNMs with this frequency only appear in the $(m_{1}+m_{2})$ GW harmonics due to the azimuthal angular selection rule. Second, different pairs of modes (up to exchange symmetry $1\leftrightarrow 2$) generally produce distinct frequencies. An exception occurs for Schwarzschild BHs, where spherical symmetry makes QNM frequencies independent of $m$. Nonetheless, different $(m_{1},m_{2})$ pairs indicate distinct coupling channels. The total QQNM amplitude in a GW harmonic is the sum of all possible pairs $(m_{1},m_{2})$ with $m_{1}+m_{2}$ fixed.
The QQNM frequency in Eq. (6) is generated by four linear QNMs:
@@
with respective strain amplitudes $A_{1}$, $A_{-1}$, $A_{2}$, and $A_{-2}$. It is worth noting that while the frequency values of $(\ell_{1},m_{1},n_{1},p_{1})$ and $(\ell_{1},-m_{1},n_{1},-p_{1})$ are related via Eq. (1), their amplitudes $A_{\pm 1}$ are independent, except in nonprecessing systems, where reflection symmetry imposes
@@
The four linear QNMs, coupled through the following four channels give rise to the same QQNM frequency:
@@
Below we will refer to them as +++++ +, ++-+ -, +-+- +, and --- -, respectively. For example, the (+)(+-)( + - ) channel corresponds to the interaction where the linear mode $(\ell_{2},-m_{2},n_{2},-p_{2})$ generates a frequency component $-\bar{\omega}_{\ell_{2}-m_{2}n_{2}}^{-p_{2}}$ in the linearized metric, which is numerically equal to $\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}}$ due to Eq. (1). This component then interacts with the frequency $\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}$ generated by the first linear mode $(\ell_{1},m_{1},n_{1},p_{1})$. Denoting the strain amplitude of the QQNM as $A_{Q}$ (using either the spherical or spheroidal basis), we have
@@
where the coefficients $R$’s are fully determined by the dimensionless spin of a Kerr BH, and the angular mode under consideration. Here we use spin-weighted spherical/spheroidal harmonics as the mode basis to facilitate direct comparison with previous results, such as those from numerical relativity and from Paper I. Alternatively, modes can also be classified by parity, see the Supplementary Material for more details.
@@
The structure of $S_{4}^{(2)}$ explicitly illustrates how two parent modes couple. The first parent mode contributes the three Weyl scalars $\Psi_{4}^{(1)}$, $\Psi_{3}^{(1)}$, and $\Psi_{2}^{(1)}$, which are then multiplied by the terms contributed by the second parent mode (indicated in parentheses). For a given linear QNM $(l,m,n,p)$, the expressions for $\Psi_{4}^{(1)}$, $\Psi_{3}^{(1)}$, and $\Psi_{2}^{(1)}$ were analytically derived in Paper I [18], see Eqs. (38), (42), and (46) therein. An intriguing feature emerges: all three Weyl scalars contain only a single frequency component, $\omega_{lmn}^{p}$, with no contribution from the $-\bar{\omega}_{lmn}^{p}$ frequency component in the linearized metric that is associated with the mode $(\ell,m,n,p)$. As a result, the source in Eq. (12) contains no terms resulting from the product of two conjugated frequency components.
@@ IV Results and Comparisons
The QQNM $\begin{pmatrix}220+\\ 220+\end{pmatrix}$ in the $l=m=4$ harmonic has been extensively discussed in the literature [12, 10, 11, 13, 14, 19, 20, 21, 18]. In particular, in Paper I, we found that the computed amplitude ratio, $R_{++}$ for the $(4,4)$ mode in Eq. (10), was smaller than those extracted from numerical relativity [13, 14]. We speculated that the difference was due to parity freedom in the parent modes (see Sec. VI. B in [18]). This was later confirmed in [19, 20, 21] for Schwarzschild BHs. In this work, after classifying all mode-coupling channels, we revisit the discrepancy for Kerr BHs.
@@
Besides $\begin{pmatrix}220+\\ 220+\end{pmatrix}$, we also study six more QQNMs, including $\begin{pmatrix}330+\\ 220+\end{pmatrix}$, $\begin{pmatrix}440+\\ 2\text{-}20-\end{pmatrix}$, $\begin{pmatrix}440+\\ 3\text{-}20-\end{pmatrix}$, $\begin{pmatrix}220+\\ 200+\end{pmatrix}$, $\begin{pmatrix}220+\\ 200-\end{pmatrix}$, and $\begin{pmatrix}330+\\ 2\text{-}20-\end{pmatrix}$. These modes are heuristically picked, motivated by the observational relevance of the $(3,3,0,+)$ mode [^2] Other linear modes, such as $(210+)$ [29, 46], are also relevant. and the the recent QQNMs discovered in numerical waveforms [47]. We use polynomials to fit their $R$ for different channels as functions of $\chi$:
@@
V Implications for future ringdown data analysis
We now analyze the significance of the QQNMs on future ringdown data analysis by including all channels. For a binary system, the remnant surrogate model from [48] predicts the remnant’s spin, which can be used in Eq. (13) to compute quadratic amplitude ratios. When combined with the postmerger ringdown model [8], these ratios enable precise quadratic amplitude predictions. This information has the potential to improve future Bayesian analyses by narrowing prior ranges and reducing the parameter space. Note that our calculations include precessing systems.
Designing a systematic Bayesian framework that incorporates our results is beyond the scope of this paper. Instead, as a figure of merit, we illustrate a practical application of our calculations by estimating the SNR of various QQNMs that have not been previously discussed in the literature. Below, we follow the method outlined in [49, 50, 43] to compute these SNRs. For simplicity, we average the SNRs over all angles, which yields the orthogonal condition between two antenna patterns $F_{+,\times}$ for CE:
@@ V Implications for future ringdown data analysis
For LISA, there is an additional factor of $3/2$. As pointed out in Paper I, a QQNM is present in different $\ell$ harmonics. Its total SNR is given by
@@
Figure 2: The SNR dependence of the QQNM, $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$, on progenitor binaries’ mass ratio $q$ and effective spin $\chi_{\rm eff}$, assuming CE. The binary systems have a total mass of ${60}M_{\odot}$ and a luminosity distance of 410 Mpc.
@@
In this paper, we adopt a conservative estimate by choosing the starting time to be $10M_{\rm tot}$ after the peak, where $M_{\rm tot}$ is the total mass of the binary system. Selecting an earlier starting time can moderately increase the SNRs, since QQNMs decay faster than linear fundamental modes.
Figure 2 shows the SNR dependence of the QQNM $\begin{pmatrix}330+\\ 220+\end{pmatrix}$ on the mass ratio $q$ and the effective inspiral spin $\chi_{\rm eff}$. This analysis is performed for CE using $M_{\rm tot}=60M_{\odot}$ and a luminosity distance of 410 Mpc. For simplicity, we set both component spins equal, $\chi_{1}=\chi_{2}=\chi_{\rm eff}$. We can see that the optimal binary configuration occurs at $q\sim 2.75$ and maximally positive $\chi_{\rm eff}$, yielding SNR $\sim 4.7$.
Here, we focus on a representative binary by fixing $M_{\rm tot}$ and $D_{L}$. The total mass affects the SNR in two ways: (1) it scales the overall strain linearly; and (2) it sets the central frequency of a QQNM. Varying $M_{\rm tot}\in[20,200]M_{\odot}$, we find that SNR increases monotonically from 0.2 to 59. The corresponding plot is included in the Supplementary Material.
Besides $\begin{pmatrix}330+\\ 220+\end{pmatrix}$, Table 1 summarizes the SNRs of various QQNMs. Most are significant enough for LISA detection, while for CE, the SNRs fall just below the typical threshold ($\sim 8$) at the given $M_{\rm tot}$ and $D_{L}$, though they could be higher for more massive or closer binaries. Nevertheless, our results suggest the importance of subdominant QQNMs and the difficulty of future ringdown data analysis: While the QQNMs may lack sufficient SNR for precise parameter estimation, omitting them — as noted in [47] — can introduce systematic errors in linear mode fitting, destabilizing extracted amplitudes. Thus, a robust ringdown analysis pipeline will be essential for future detectors.
Table 1: Optimal SNRs of various QQNMs, assuming CE with $M_{\rm tot}=60M_{\odot}$ and $D_{L}=410$ Mpc, and LISA with $M_{\rm tot}=8\times 10^{6}M_{\odot}$ and a redshift of 1. The SNRs are obtained from optimal binary configurations that maximize SNR for the specified $M_{\rm tot}$ and $D_{L}$.
QQNM Optimal SNR
configuration CE LISA
$\begin{pmatrix}2,&2,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$ $q\sim 1$ large $\chi_{\rm eff}$ 9 970
$\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$ $q\sim 2.75$ large $\chi_{\rm eff}$ 4.7 653
$\begin{pmatrix}4,&4,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$ $q\sim 4.25$ large $\chi_{\rm eff}$ 3 140
$\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$ $q\sim 2.75$ large $\chi_{\rm eff}$ 5 60
$\begin{pmatrix}4,&4,&0,&+\\ 3,&-2,&0,&-\end{pmatrix}$ $q\sim 8$ large $\chi_{\rm eff}$ $0.084$ 2.1